OpenStudy (anonymous):

how do I solve for this factored equation using zero property? -4(4x+1)(x-2): 4x+1=0 x=-1/4 and x=2 but what about the first -4?

OpenStudy (anonymous):

The first -4 makes no difference if either (4x+1) =0 or (x-2) =0 because if multiplied by 0 it equals 0.

OpenStudy (anonymous):

so answer is -1/4 and 2?

OpenStudy (whpalmer4):

\[-4(4x+1)(x-2) = 0\]We can divide both sides by -4 giving us \[(4x+1)(x-2) = 0\]Now for that to be true, we need to find all values of x that make either product term be 0, so \[4x+1=0\]\[4x=-1\]\[x=-1/4\]\[x-2=0\]\[x=2\] You got the right answers!